Q1.Calculate the current I by superposition. ЧА Чл 128 3л 6л. I дох 55L
Answer #1

Superposition Theorem It states that for a leneul bilateral electrical circuit with more than one the gesponse auty element SCalculating I, and Now we have remove all sounes except 28 V source Heplace them with internal rogistance be, current sourceVw [*+ 4+ 1 - 135 - 80 50 64 Vx/13 12 +80 = 64 +80 11 644 840 3 13 X9 = 304 Vx = 4x304 13 1216 13 Vx = 1216 13 from circuit dYA VB 3.2 Vc. Here in VA-VB VA w it 4.2 dlx 4 $6.2 zoia 5.1 2.12 By applying kch at VA node VA 2 VA + ve +4 70 Va[+6] +vo FH3 VA-VB Ve-VB + V-30 t 3 2 5 Vc - VB VA-VB Vc + c els = V-ve) 3 2 = 4 5 VEET 3 + + + 5+ V8 (5) 3 IS VA-VB) = 4 (2x8) Ve ( 1 8Substitution en Cariation 3): 1:1) 69 VB + V3 (24) tve ( 32) 20 16 - VB 3x4 + + VB (3) + VC C co 3x4 12-#) + ( VBC +Vc (-1/2)Ne (ch 13 15 11 12-4 3 culos Vc = -8x15 13X3 --120 39 Vea -120 39 :-3077 V Vc =-3.0920 By substituting ve cen qreation : Ve +Ig= VB from circuit diagram: - 3.538 6 Ig=-0.5896 A from circuit diagram: Ig = Ve -92/26 -92 il I! -92 26 X6 = 6 156 Ig - -92